The probability (i) that A can solve a problem in statistics in 4/5 (ii) that B can solve it in 2/3 (iii) that C can solve it in 3/7, if all them try independently. Find the probability that the problem will be solved
1
Expert's answer
2020-04-06T17:23:23-0400

The problem will not be solved when A,B, and C fail to solve.

Probability of all options failing to solve is "\frac{1}{5}×\frac{1}{3}×\frac{4}{7}=\frac{4}{105}"

Thus, the probability of the problem being solved is "1-\frac{4}{105}=\frac{101}{105}"